OpenStudy (dariyana):

Using complete sentences, explain what might happen if the order of operations was used to solve a literal equation.

OpenStudy (dariyana):

\[\frac{ S }{ 4 }=\frac{ 4V^{2} }{ 4? }\]

ganeshie8 (ganeshie8):

you should not use order of operating when solving a literal equation. it complicates the isolation process

ganeshie8 (ganeshie8):

for example, take below equation and try to solve using 'order of operations' :- Solve the literal equation : 2x-3 = 7

ganeshie8 (ganeshie8):

PEMDAS first comes Multiplication, since 2 is attached to x, to isolate x, we divide the whole equation with 2 : 2x-3 = 7 /2 /2 x - 3/2 = 7/2

ganeshie8 (ganeshie8):

Next, add 3/2 both sides x - 3/2 = 7/2 + 3/2 +3/2

ganeshie8 (ganeshie8):

x = 7/2 + 3/3 x = 10/2 x = 5

ganeshie8 (ganeshie8):

Obviously this is a bit complicated. You sure know the the other easy way to solve this... dont you ?